# Beyond the Main Sequence: Binary Evolution Pathways to Blue Stragglers in the Gaia Era

### v1

@@ B.2 Isochrone-fitting models
Previous studies have attempted to extract BSS parameters based on isochrone models such as JS21 and D23. In the first study the authors assume the mass of the BSSs to be approximately, the mass of a point at the ZAMS with the same magnitude. In the second study, they fit multiple isochrones with different ages and compare BSS positions with respect to isochrones. We take an approach similar to the second one. For GCs (OCs), we use multiple PARSEC isochrones (Bressan et al. 2012) from 200 Myr (100 Myr) in steps of 50 Myr (25 Myr) up to the cluster age found with the best-isochrone fitting, fixing the metallicity from the literature for each cluster as described earlier. We select the three closest points on the best-fit isochrone for each cluster and compute the average and rms for mass, luminosity, age, surface gravity and effective temperature for each BSS.
@@ B.3 Spectroscopy from Gaia DR3
Gaia DR3 includes optical low–resolution spectra for over 220 million stars. As a brief summary, these spectra are measured with two instruments. One covering the blue–optical range (380–680 nm, measured with the “Blue Photometer” instrument) called BP and another covering infrared range (640–1050 nm, measured with the “Red Photometer”) called RP (De Angeli et al. 2023; Montegriffo et al. 2023). These BP/RP spectra, hereafter referred to as XP spectra, have a spectral resolution with parameters inferred from “classical” models (Andrae et al. 2023, and references therein), that are highly sensitive to errors in low-resolution spectra, which is the case for Gaia DR3 XP spectra. In another approach, Zhang et al. (2023), based on parameters obtained from higher-resolution spectra () from other surveys, cross-matched them with the Gaia DR3 spectra and created a machine-learning model that has the advantage of: i) obtaining astrophysical parameters from higher-resolution spectra, compare them with the XP spectra and adjust those parameters at lower spectral resolution; ii) being able to find various systematic errors that could not easily be noticed in the XP spectra alone. For this reason, we extract XP parameters (hereafter called XPP) from Zhang et al. (2023) for all our sample BSSs, if available, such as effective temperature and surface gravity. We cross-match the provided parameters with our selected BSSs using Gaia Source ID unique identifier. Finally, Zhang et al. (2023) provides a quality_flag parameter which, in short, is a number to quantify the reliability of the obtained parameters from spectra. The authors suggest a “basic reliability cut” to be at least (note, a lower number indicates more reliable parameters). We decided to only use stars with to ensure we extract only reliable results. More details for these parameters are provided in Section 4 of Zhang et al. (2023).
@@
From the IRFM we find for GCs and for OCs. From isochrone fitting we find for GCs and for OCs. Only 537 out of 4399 BSSs ( of our sample) had XPP parameters measured for 279 stars in GCs and 258 stars in OCs. With this method, we find the mean BSS effective temperature of for GCs and for OCs. Figure 13 shows the distribution for , , and of all our sample BSSs. We have also estimated the parameters for YSSs and RSSs which will be available online as electronic tables. Their distributions are shown in Figure 15 and Figure 16 respectively.
@@
To check the consistency of our results, we compare the values obtained from the three methods in Figure 14 against each other. The top figure is divided into two sub-panels, both of which are comparing derived with the isochrone fitting method against the values found with the IRFM (top–upper panel) and XPP (top–lower panel). The lower panel compares the values of latter two methods against each other. We only consider BSSs with an error in of less than in both components. The 1–to–1 relation is shown as a gray dashed line and compared with the data. We observe very good agreement between the methods, albeit with small level of systematics. We fit a linear model using LMFIT Python package777https://lmfit.github.io/lmfit-py/ to these relations considering full error propagation in all values, which are shown as a dot-dashed green lines. We also estimated pearson r-coefficients for all these relations in a monte carlo approach, and the estimated values are 0.93 (top panel), 0.76 (middle panel) and 0.74 (bottom panel). The errors correspond to 16th and 84th percentile values with respect to the median value. The r-coefficients suggest that estimated from different methods are strongly correlated. We obtain the following linear relations in Kelvin units:
@@
From these results, we note that extracted from isochrone fitting and those derived from the color–temperature relations (IRFM) appear most consistent with each other up to , with little systematic offset towards higher IRFM temperature. A similar behaviour is observed for the XPP vs. isochrone method, where values are consistent within the scatter up to . Above that temperature, a small systematic offset towards lower XPP temperatures is observed. The bottom panel of Figure 14 shows the XPP vs. IRFM method comparison, which shows the largest scatter yet fairly consistent values for most BSSs, while the linear fit is pulled from 1–to–1 relation by some OC BSSs with temperatures above .
@@
From these comparisons, we estimate that the systematic uncertainty between various methods is of the order K. The IRFM and isochrone methods consider the metallicity of the cluster, while the XPP method is inferior for the reasons described in Section B.3, which is reflected in the lower scatter and smaller systematic offset in the IRFM vs. isochrone method comparison (upper panel in Fig. 14). In addition, XPP parameters are not available for stars at the MSTO and we, therefore, are not able to extract using XPP. We, therefore, consider the IRFM and isochrone method from now as the bona-fide best estimates, keeping in mind the systematics described above.
@@ Appendix C Astrophysical parameters for YSSs and RSSs
Similar to Figure 13, we display the extracted astrophysical parameters for stars classified as Yellow Straggler Stars (YSSs) and Red Straggler Stars (RSSs) in Figures 15 and 16, respectively.

### v2

@@ B.2 Isochrone-fitting models
Previous studies have attempted to extract BSS parameters based on isochrone models such as JS21 and D23. In the first study, the authors assumed that the mass of the BSSs to be approximately the mass of a point at the ZAMS with the same magnitude. In the second study, they fit multiple isochrones with different ages and compare BSS positions with respect to the isochrones. We take an approach similar to the second one. For GCs (OCs), we use multiple PARSEC isochrones (Bressan et al. 2012) from 200 Myr (100 Myr) in steps of 50 Myr (25 Myr) up to the cluster age found with the best-isochrone fitting, fixing the metallicity from the literature for each cluster as described earlier. We select the three closest points on the best-fit isochrone for each cluster and compute the average and rms for mass, luminosity, age, surface gravity and effective temperature for each BSS.
@@ B.3 Spectroscopy from Gaia DR3
Gaia DR3 includes optical low–resolution spectra for over 220 million stars. As a brief summary, these spectra are measured with two instruments. One covers the blue–optical range (380–680 nm, measured with the “Blue Photometer” instrument) called BP and another covers the infrared range (640–1050 nm, measured with the “Red Photometer”) called RP (De Angeli et al. 2023; Montegriffo et al. 2023). These BP/RP spectra, hereafter referred to as XP spectra, have a spectral resolution with parameters inferred from “classical” models (Andrae et al. 2023, and references therein), which are highly sensitive to errors in low-resolution spectra, which is the case for Gaia DR3 XP spectra. In another approach, Zhang et al. (2023), based on parameters obtained from higher-resolution spectra () from other surveys, cross-matched them with Gaia DR3 spectra and created a machine-learning model that has the advantage of: i) obtaining astrophysical parameters from higher-resolution spectra, comparing them with the XP spectra and adjust those parameters at lower spectral resolution; ii) being able to find various systematic errors that could not easily be noticed in the XP spectra alone. For this reason, we extract XP parameters (hereafter called XPP) from Zhang et al. (2023) for all our sample BSSs, if available, such as effective temperature and surface gravity. We cross-match the provided parameters with our selected BSSs using Gaia Source ID unique identifier. Finally, Zhang et al. (2023) provides a quality_flag parameter which, in short, is a number to quantify the reliability of the obtained parameters from spectra. The authors suggest a “basic reliability cut” to be at least (note, a lower number indicates more reliable parameters). We decided to only use stars with to ensure we extract only reliable results. More details for these parameters are provided in Section 4 of Zhang et al. (2023).
@@
From the IRFM we find for GCs and for OCs. From isochrone fitting we find for GCs and for OCs. Only 537 of 4399 BSSs ( of our sample) had XPP parameters measured for 279 stars in GCs and 258 stars in OCs. With this method, we find the mean BSS effective temperature of for GCs and for OCs. Figure 14 shows the distribution for , , and of all our sample BSSs. We also estimated the parameters for YSSs and RSSs, which are available at the CDS. Their distributions are shown in Figure 16 and Figure 17 respectively.
@@
To check the consistency of our results, we compare the values obtained from the three methods in Figure 15 against each other. The top figure is divided into two sub-panels, both of which are comparing derived with the isochrone fitting method against the values found with the IRFM (top–upper panel) and XPP (top–lower panel). The lower panel compares the values of latter two methods against each other. We only consider BSSs with an error in of less than in both components. The 1–to–1 relation is shown as a gray dashed line and is compared with the data. We observe very good agreement between the methods, albeit with small level of systematics. We fit a linear model using LMFIT Python package666https://lmfit.github.io/lmfit-py/ to these relations considering full error propagation in all values, which are shown as a dot-dashed green lines. We also estimated pearson r-coefficients for all these relations in a monte carlo approach, and the estimated values are 0.93 (top panel), 0.76 (middle panel) and 0.74 (bottom panel). The errors correspond to 16th and 84th percentile values with respect to the median value. The r-coefficients suggest that estimated from different methods are strongly correlated. We obtain the following linear relations in Kelvin units:
@@
From these results, we note that the extracted from isochrone fitting and those derived from the color–temperature relations (IRFM) appear most consistent with each other up to , with little systematic offset towards higher IRFM temperature. A similar behavior is observed for the XPP versus isochrone method, where values are consistent within the scatter up to . Above that temperature, a small systematic offset towards lower XPP temperatures is observed. The bottom panel of Figure 15 shows the XPP versus IRFM method comparison, which shows the largest scatter yet fairly consistent values for most BSSs, while the linear fit is pulled from 1–to–1 relation by some OC BSSs with temperatures above .
@@
From these comparisons, we estimate that the systematic uncertainty between various methods is of the order K. The IRFM and isochrone methods consider the metallicity of the cluster, while the XPP method is inferior for the reasons described in Section B.3, which is reflected in the lower scatter and smaller systematic offset in the IRFM versus isochrone method comparison (upper panel in Fig. 15). In addition, XPP parameters are not available for stars at the MSTO and therefore we are not able to extract using XPP. We therefore consider the IRFM and isochrone method from now as the bona fide best estimates, keeping in mind the systematics described above.
@@ Appendix C Astrophysical parameters for YSSs and RSSs
Similarly to Figure 14, we display the extracted astrophysical parameters for stars classified as Yellow Straggler Stars (YSSs) and Red Straggler Stars (RSSs) in Figures 16 and 17, respectively.
Appendix D Parameters of star clusters studied in this paper and the properties of BSSs detected in those clusters
cluster cluster RA Dec log(age/yr) Av distance [Fe/H] Teff,IRFM Teff,isochrone NBSS NMS fBSS/MS name type (deg) (deg) (mag) (pc) (dex) (K) (K) (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) NGC104 GC 6.175 72.024 10.0700.011 0.124 4367 0.65 590392 5995 415 13691 0.060.002 NGC288 GC 13.381 26.529 10.0640.002 0.031 9000 0.8 6066116 6173 150 – – NGC362 GC 15.885 70.781 10.0680.013 0.155 9090 1.3 6248160 6368 137 1578 0.1710.011 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ Berkeley39 OC 116.893 4.735 9.811 0.029 0.394 4156 0.0 5932 84 6038 22 264 0.14 0.025 Berkeley32 OC 104.738 6.379 9.686 0.022 0.553 3321 0.3 6286 125 6403 14 139 0.151 0.035 Collinder261 OC 189.802 68.492 9.898 0.065 0.999 2706 0.1 5772 72 5865 64 460 0.228 0.025 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ 777The complete table is made available at the CDS. The list of columns represented here is: (1) Name of the cluster; (2) whether it is a globular cluster (GC) or open cluster (OC); (3) and (4) represents central coordinates of the cluster, (5) age of the cluster in log scale, (6) extinction (AV) towards the cluster, (7) distance to the cluster, (8) metallicity of the cluster, (9) mean effective temperature of BSS populations in that cluster using IRFM method, (10) mean effective temperature of BSS populations in that cluster using isochrone method; (11) number of BSSs detected in that cluster; (12) number of main sequence (MS) stars in the luminosity range between the MS turn-off (MSTO) and MSTO+1 mag in that cluster; (13) fractional BSSs population to MS populations.
isochrone method cluster type Gaia source id Teff,IRFM Teff log(age/yr) M log(L/L⊙) logg Teff,xpp loggxpp Me Mesequence (K) (K) (M⊙) (K) (mag) (mag) (1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11) (12) (13) (14) (15) NGC2141 OC 3341815582307132416 636382 651059 9.2220.015 1.700.01 1.0400.006 3.8330.010 – – – – 0.17 lower 2.237 0.597 NGC2141 OC 3341815582307133056 639184 646479 9.2220.015 1.720.00 1.0700.003 3.8000.020 – – – – 0.19 lower 2.176 0.589 NGC2141 OC 3341815681086838272 6736101 685595 9.0870.077 1.560.01 0.8000.005 4.1220.025 – – – – 0.08 lower 2.807 0.500 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ NGC104 GC 4689572883825672192 6422126 656746 9.7400.027 1.010.01 0.5500.004 4.1190.015 6831136 4.1420.123 0.28 upper 3.538 0.566 NGC104 GC 4689573433581464192 6262109 638426 9.8420.024 0.930.01 0.3500.003 4.2260.009 6896166 4.2160.2 0.18 lower 4.043 0.609 NGC104 GC 4689574086416430464 6669154 683835 9.5050.035 1.080.01 0.4900.002 4.2710.009 6652159 4.1950.142 0.37 upper 3.651 0.505 ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ ⋮ 888The complete catalog is made available at the CDS. The list of columns presented here is (1) the name of the cluster to which the BSS belong; (2) whether the cluster is classified as an open cluster (OC) or globular cluster (GC); (3) source id of the BSS as per the Gaia DR3; (4) effective temperature of the BSS based on IRFM method; (5) – (9) effective temperature, age, mass, luminosity and surface gravity of the BSS, respectively, estimated using isochrone method; (10) – (11) effective temperature and of the BSS estimated from Gaia-XPP spectrum. (12) fractional mass excess of the BSS based compared to MSTO mass; (13) location of the BSS based on mass excess, as mentioned in Figure 9; (14)-(15) BP magnitude and (BP-RP) color in absolute scale.